# DRUPI: Dataset Reduction Using Privileged Information

### v2

@@ C.2 Further Regression Supervision Objectives
Table 12: Ablation study of different components in DRUPI across CIFAR-10 and CIFAR-100 (1 IPC). The table compares the effects of feature labels, CE loss ($\mathcal{L}_{CE}$), contrastive loss ($\mathcal{L}_{IN}$), and soft labels on model performance.
Feature label $\mathcal{L}_{CE}$ $\mathcal{L}_{IN}$ Soft label Acc (%)
@@
(a) CIFAR-10, 1 IPC
Feature label $\mathcal{L}_{CE}$ $\mathcal{L}_{IN}$ Soft label Acc (%)
@@
(b) CIFAR-100, 1 IPC
Appendix D Pseudo Code of DRUPI
Algorithm  1 outlines the process where we initialize the reduced dataset with assigned feature labels based on the synthetic dataset from dataset condensation (DC). Subsequently, we update the reduced dataset by learning the feature labels, which are progressively refined during the training process.
Algorithm 1 Dataset Reduction Using Privileged Information (DRUPI) For DC

### v3

@@ C.2 Further Regression Supervision Objectives
Table 13: Ablation of DCPI components on CIFAR-10 with IPC = 1 and CIFAR-100 with IPC = 1. The table compares the effects of feature labels, CE loss ($\mathcal{L}_{CE}$), contrastive loss ($\mathcal{L}_{IN}$), and soft labels on model performance.
Feature $\mathcal{L}_{CE}$ $\mathcal{L}_{IN}$ Soft Acc (%)
@@
(a) CIFAR-10, fraction = 0.02%
Feature $\mathcal{L}_{CE}$ $\mathcal{L}_{IN}$ Soft Acc (%)
@@
(b) CIFAR-100, fraction = 0.2%
Appendix D Pseudo Code of DCPI
Algorithm 1 outlines the process where we initialize the reduced dataset with assigned feature labels based on the synthetic dataset from dataset condensation (DC). Subsequently, we update the reduced dataset by learning the feature labels, which are progressively refined during the training process.
Algorithm 1 Dataset Condensation using Privileged Information (DCPI) For DC
@@ Appendix D Pseudo Code of DCPI
D.1 Theoretical Analysis
We analyze how the DCPI improves reduced dataset quality. Let $g_{\mathcal{T}}\in\mathcal{G}_{\mathcal{T}}$ be the oracle function for dataset $\mathcal{D}_{\mathcal{T}}$, and $|\cdot|_{C}$ denote a model performance measure. Consider two models: $g_{\mathcal{S}}\in\mathcal{G}_{\mathcal{S}}$ trained on reduced dataset $\mathcal{D}_{\mathcal{S}}$, and $g_{\mathcal{S}^{\star}}\in\mathcal{G}_{\mathcal{S}^{\star}}$ trained on $\mathcal{D}_{\mathcal{S}^{\star}}$ with privileged information.
Using VC theory Vapnik (1998), we assess its performance via VC-dimension. For a model $g\in\mathcal{G}$ with finite VC-dimension $|\mathcal{G}|_{\text{VC}}$, the expected error $R(g)$ is bounded with probability $1-\delta$ as:
where $O(\cdot)$ is the estimation error, $R_{m}(g)$ is the training error over $m$ data points, and $\alpha\in[\frac{1}{2},1]$ reflects task difficulty. For non-separable tasks, $\alpha\approx\frac{1}{2}$, giving a learning rate of $O(m^{-1/2})$. For separable tasks with no training errors, $\alpha\approx 1$, yielding $O(m^{-1})$. A teacher model can accelerate learning from $O(m^{-1/2})$ to $O(m^{-1})$ for a student on a dataset of size $m$.
We analyze DCPI, showing how privileged information speeds up learning and enhances reduced dataset quality. Extending Lopez-Paz et al. (2015), we apply results to dataset condensation. The model $g_{\mathcal{S}}$ on reduced dataset $\mathcal{D}_{\mathcal{S}}$ learns the true function $g_{\mathcal{T}}$ at a slower rate $\alpha_{\mathcal{S}}$:
where the final inequality arises because $\alpha\leq 1$. Therefore, for $\frac{1}{2}<\alpha\leq 1$ in dataset condensation settings, the inequality becomes:
This inequality highlights the advantages of models trained with privileged information: training using privileged information exhibit lower generalization and approximation errors compared to those trained without privileged information. More importantly, it emphasizes that the privileged information is most beneficial in low-data regimes, which is the typical DR scenario. These benefits align with the principles of LUPI as outlined in  Vapnik et al. (2015); Lopez-Paz et al. (2015).
Appendix E Further Discussion about DCPI
The advancement of dataset condensation techniques like DCPI holds significant implications for the broader machine learning landscape. By enabling the creation of smaller, yet more statistically potent datasets, DCPI can democratize access to training sophisticated models, particularly for entities with constrained computational resources. This could accelerate research and development in various domains by lowering the barrier to entry for working with large-scale data. The ability to synthesize richer supervisory signals beyond simple data-label pairs may also foster the development of more data-efficient learning algorithms and contribute to models with enhanced generalization capabilities. From a sustainability perspective, training on compact datasets naturally translates to reduced energy consumption and a smaller carbon footprint associated with model development cycles. However, it is crucial to consider the potential for bias amplification. If the original large dataset, or the teacher models used implicitly or explicitly in the synthesis of privileged information, contain inherent biases, DCPI could inadvertently concentrate or even reshape these biases within the reduced dataset. Therefore, rigorous auditing of both the source data and the characteristics of the synthesized privileged information will be essential to ensure that the benefits of dataset condensation do not come at the cost of fairness or equity in downstream applications.